OpenStudy (alivejeremy):

Solve. 4y + 16 > –8 A. y > –6 B. y > –2 C. y > 2 D. y > 6

OpenStudy (anonymous):

OK, first pretend that the > sign is and equal.

OpenStudy (anonymous):

Just remember if your y is negative, it will swap the sign.

OpenStudy (alivejeremy):

yea

OpenStudy (anonymous):

So now you have 4y+16=-8 Solve for y

OpenStudy (alivejeremy):

is it 6

OpenStudy (alivejeremy):

or 20

OpenStudy (anonymous):

Almost.

OpenStudy (alivejeremy):

-6

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

Now put the > sign replacing the = so y=-6

OpenStudy (alivejeremy):

si the answer is A

OpenStudy (alivejeremy):

so*

OpenStudy (anonymous):

yes

OpenStudy (alivejeremy):

?

OpenStudy (anonymous):

Do you know what \[\le \ge\] mean?

OpenStudy (alivejeremy):

yea wbu

OpenStudy (anonymous):

Follow the direction or the red arrow.

OpenStudy (anonymous):

*of

OpenStudy (alivejeremy):

?

OpenStudy (alivejeremy):

oh

OpenStudy (anonymous):

Which means your answer is...

OpenStudy (alivejeremy):

C

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

More?

OpenStudy (alivejeremy):

^_^

OpenStudy (anonymous):

First, take the 3 to the other side.

OpenStudy (alivejeremy):

3?

OpenStudy (anonymous):

When you take something to the opossite you do the opossite.

OpenStudy (anonymous):

So the opossite of division is ...

OpenStudy (alivejeremy):

multiplication

OpenStudy (anonymous):

So take the 3 to the other side you get

OpenStudy (anonymous):

the answer for this is  http://assets.openstudy.com/updates/attachments/567aa50ae4b0f8b8229fa56b-alivejeremy-1450878890814-capture.png

OpenStudy (alivejeremy):

\[\frac{ 3 }{s }\]

OpenStudy (anonymous):

@siblings Don't give direct answers. It's against the code of conduction.